A car covers its outward journey at 40kmph. It is towed back, over exactly the same distance at 10mph. What is the average speed over the entire journey?
 
A)12mph B)14mph C)15mph D)20mph
 
 
Ans :D

Grade:10

2 Answers

Arun
6 years ago
Dear student
 
we can apply the formula 2xy/(x+y) for average speed.
Hence
 
Avg speed = 2*40*10/(40+10) = 800/50 = 16 kmph
 
Regards
Arun (askIITians forum expert)
Arun
6 years ago
 
Sorry,the above soltuion is incorrect due to not taking conversion
 
40 kmph = 40/1.6 mph = 24.85 mph
 
Now avg speed = 2*24.85*10/34.85 = 14.26 mph
 
Hence option B is correct